# ACTIVA: Amortized Causal Effect Estimation without Graphs via Transformer-based Variational Autoencoder

### v1

@@ 3 Model Architecture
Figure 2: Architecture of our conditional $\beta$-VAE model. Orange components are part of the encoder, while green components are part of the decoder. The dotted lines indicate the batch dimensions for MHSA. First, the dataset $\boldsymbol{D}^{\mathcal{M}}$ is concatenated with the intervention representation $\boldsymbol{I}_{i,t}$, repeated $N$ times. The resulting 3D tensor is then fed through $L$ layers of MHSA, alternating between attending over the $d$ and $N$ axis. The parameters for the posterior distribution are predicted by mean pooling over the sample axis and applying a row-wise MLP. After reparametrization, $\boldsymbol{I}_{i,t}$ is appended to $\boldsymbol{z}$ and the resulting tensor fed through $K$ layers of MHSA and, finally, a row-wise MLP to predict the parameters of the estimated interventional distribution.
We propose a method for amortized prediction of interventional distributions from observational data using a conditional $\beta$-VAE with VAMP prior. Our model employs a transformer-based encoder (Kossen et al., 2021; Lorch et al., 2022) and decoder that predicts the parameters of a mixture of Gaussians which represents the interventional distribution. This formulation allows us to model arbitrary interventional distributions that are permutation equivariant regarding the variables, i.e. permuting the variable ordering in the input data leads to an equivalent permutation of the dimension of the estimated distribution. Figure 2 gives an overview of our model. More details and hyperparameters can be found in Appendix C.
@@ 3.1 Representing Interventions
To effectively condition our VAE on a given intervention, we create a matrix representation of the interventions. More precisely, we consider $i\in\{1,\dots,|I|\}$ the index of a possible intervention value $v_{i}$, where $|I|$ is the number of possible values, and a selector $\boldsymbol{t}\in\{0,1\}^{d}$ indicating the target(s) of the intervention. By not encoding the intervention value directly, but rather the index of the intervention, we remove the need to have a full specification of the intervention. This allows us to model interventional distributions based on placeholder IDs of the interventions, as long as they can be attributed to interventional training data.
@@
From $i$ and $\boldsymbol{t}$ we construct the intervention representation $\boldsymbol{I}_{i,t}$ representing $do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i})$ as follows. We perform a one-hot encoding of $i$ creating a vector $\boldsymbol{i_{oh}}\in\mathbb{R}^{\mid I\mid}$ and repeat it $d$ times, resulting in a matrix $\boldsymbol{i_{rep}}\in\mathbb{R}^{d\times\mid I\mid}$. We then apply $\boldsymbol{t}$ as a mask to this matrix, effectively zeroing out the rows that correspond to non-intervened variables.
@@
This construction of the conditional of our VAE ensures that the intervention-relevant information for each variable is provided as local information alongside the variable itself. This design guarantees that the information for one variable is stored independently of all other variables, ensuring permutation equivariance of the model with respect to the variable ordering in the data even after conditioning on the intervention.
3.2 Encoder
@@
We model our encoder based on the extension of non-parametric encoders (Kossen et al., 2021) described in (Lorch et al., 2022). As it has been shown for various causal tasks, this encoder architecture can successfully encode a dataset into a vector containing relevant information about the underlying causal model, such as causal structure (Lorch et al., 2022), topological ordering (Scetbon et al., 2024), informative interventions (Annadani et al., 2025), or more abstract prediction tasks (Hollmann et al., 2025). In our setup, we employ this architecture to encode information necessary to predict interventional distributions.
@@
Given a dataset $\boldsymbol{D}^{\mathcal{M}}\in\mathbb{R}^{N\times d}$ of $N$ observational samples from $\mathcal{M}$, an intervention index $i$ and a binary target vector $\boldsymbol{t}$, the encoder predicts the mean $\boldsymbol{\mu}_{enc}$ and variance $\boldsymbol{\sigma}_{enc}$ of a multivariate normal Gaussian (MVN) representing the interventional distribution. To do so, we first extend the intervention representation $\boldsymbol{I}_{i,t}$ by repeating it $N$ times. We then append the expanded representation to the dataset resulting in an augmented dataset $\boldsymbol{D}^{\mathcal{M}}_{it}\in\mathbb{R}^{N\times d\times\mid I\mid+1}$. Then we apply $L$ blocks of multi-head self-attention (MHSA) that alternate in attending over $d$ features and $N$ samples. This ensures permutational equivariance to both the sample and the feature dimension (Lorch et al., 2022). Furthermore, we apply layer normalization, residual connections, and dropout following the transformer block setup (Vaswani et al., 2017). After the transformer blocks, we average over the sample axis to obtain an embedding $\boldsymbol{h}_{dec}\in\mathbb{R}^{d\times e}$, where $e$ indicates the embedding dimension. By taking the average, we ensure that the embedding is permutation invariant regarding the sample dimension. To obtain the mean $\boldsymbol{\mu}_{enc}$ and log-variance $\log\boldsymbol{\sigma}_{enc}$ we apply a multi-layer perceptron (MLP) to the embedding of each feature independently. Overall, $[\boldsymbol{\mu}_{enc},\boldsymbol{\sigma}_{enc}]=enc_{\theta}(\boldsymbol{D}^{\mathcal{M}},i,\boldsymbol{t})$ characterizes a standard normal distribution, where $\theta$ are the parameters of our encoder.
@@ 3.3 Decoder
We model the decoder $dec_{\phi}(\boldsymbol{z},i,\boldsymbol{t})$ as a transformer with parameters $\phi$ that outputs parameters of a mixture of Gaussians with $c$ components, where $\boldsymbol{z}$ is an encoding, $i$ indicates a possible intervention value $v_{i}$, and $\boldsymbol{t}$ defines the target vector. This allows us to get a closed-form expression of the estimated interventional distributions.
[…]
Since $\boldsymbol{z}$ encodes causal information as argued in Section 3.2, by conditioning on an intervention $do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i})$ via $i$ and $\boldsymbol{t}$, we enable our model to estimate interventional distributions. It is noteworthy that we maintain permutation equivariance of the means and covariances of the resulting mixture w.r.t. the ordering of the variables in $\boldsymbol{z}$. This allows us to propagate the permutations of the variables in the input data, through the encoder and the decoder to the estimated interventional distribution.
3.4 Learning Objective
Our model optimizes a conditional version of the ELBO on the log-likelihood of the data. Specifically, we aim to maximize the log-likelihood of the interventional data $\log p_{\mathcal{M}}(\boldsymbol{x}^{\mathcal{M}_{do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i})}}|\boldsymbol{D}^{\mathcal{M}},do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i}))$. We point out that, while SCMs are useful for describing our approach, we do not rely on any specific formalization of causal models to derive our ELBO loss.
To derive our learning objective, we begin by factorizing our conditional distribution in terms of latent variables. Furthermore, as argued in Section 3.2, we assume that the latents $\boldsymbol{z}$ together with $i$ and $\boldsymbol{t}$ contain all information necessary to generate $\boldsymbol{V}_{do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i})}$. Hence, we take advantage of the fact that $\boldsymbol{V}_{do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i})}\perp\!\!\!\!\perp\boldsymbol{D}^{\mathcal{M}}\mid\boldsymbol{z}$, resulting in the following equality.
$\log p_{\theta}(\boldsymbol{V}$ $|\boldsymbol{D}^{\mathcal{M}},do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i}))$ (4)
[…]
We then introduce the conditional variational posterior $q_{\phi}(\boldsymbol{z}|\boldsymbol{D}^{\mathcal{M}},do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i}))$, parameterized by the output of the encoder and simplify further as shown below [^2] Note that for readability we denote the variational posterior as $q_{\phi}(\boldsymbol{z}\mid.)$:
[…]
$=\log\int p_{\theta}(\boldsymbol{V}|\boldsymbol{z},do(\boldsymbol{V}_{\boldsymbol{t}}))\frac{p(\boldsymbol{z})}{q_{\phi}(\boldsymbol{z}\mid.)}q_{\phi}(\boldsymbol{z}\mid.)\,d\boldsymbol{z}$
$=\log\displaystyle\mathop{\mathbb{E}}_{q_{\phi}(\boldsymbol{z}\mid.)}\left[p_{\theta}(\boldsymbol{V}|\boldsymbol{z},do(\boldsymbol{V}_{\boldsymbol{t}}))\frac{p(\boldsymbol{z})}{q_{\phi}(\boldsymbol{z}\mid.)}\right].$ (7)
We then apply Jensen’s inequality and rearrange the resulting term into reconstruction and divergence term leading to the ELBO.
$\log\displaystyle\mathop{\mathbb{E}}_{q_{\phi}(\boldsymbol{z}\mid.)}$ $\left[p_{\theta}(\boldsymbol{V}|\boldsymbol{z},do(\boldsymbol{V}_{\boldsymbol{t}}))\frac{p(\boldsymbol{z})}{q_{\phi}(\boldsymbol{z}\mid.)}\right]$ (8)
$\geq\displaystyle\mathop{\mathbb{E}}_{q_{\phi}(\boldsymbol{z}|.)}[\log p_{\theta}(\boldsymbol{V}|\boldsymbol{z},do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i}))$
$\quad+\log\frac{p(\boldsymbol{z})}{q_{\phi}(\boldsymbol{z}|.)}]$
$=\displaystyle\mathop{\mathbb{E}}_{q_{\phi}(\boldsymbol{z}\mid.)}\left[\log p_{\theta}(\boldsymbol{V}|\boldsymbol{z},do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i}))\right]$ (9)
$\quad-\operatorname{KL}(q_{\phi}(\boldsymbol{z}|.)\parallel p(\boldsymbol{z})),$
where the first term represents the expected log-likelihood of the data given the latent variable $\boldsymbol{z}$ and the intervention, and the second term is the KL divergence between the variational posterior $q_{\phi}(\boldsymbol{z}|\boldsymbol{D}^{\mathcal{M}},do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i}))$ and the prior $p(\boldsymbol{z})$. To make this bound tractable, we approximate the true posterior $p(\boldsymbol{z})$ with the VAMP prior $p_{\phi,\omega}(\boldsymbol{z})$.
We estimate the log-likelihood with $N$ samples per $|I|\times|T|$ interventional distributions in our training set, simultaneously ensuring that the ELBO is maximized for each intervention value and target. To enable gradient-based optimization, we use the reparameterization trick (Kingma & Welling, 2013), allowing backpropagation through the stochastic sampling process of $q_{\phi}(\boldsymbol{z}|\boldsymbol{D}^{\mathcal{M}},do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i}))$. This results in the reconstruction loss
$\mathcal{L}_{\text{rec}}=-\displaystyle\mathop{\mathbb{E}}_{\begin{subarray}{c}q_{\phi}(\boldsymbol{z}\mid.)\\ I;T;N\end{subarray}}\left[\log p_{\theta}(\boldsymbol{x}^{\mathcal{M}_{do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i})}}_{n}|\boldsymbol{z},do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i}))\right].$ (10)
The second term, the KL divergence $\mathcal{L}_{\text{KL}}$, acts as a regularizer, encouraging the variational posterior to remain close to the prior distribution. Similarly to the reconstruction loss, we compute this simultaneously for all intervention-target pairs, resulting in the KL loss
$\mathcal{L}_{\text{KL}}=\mathbb{E}_{I;T}\left[\operatorname{KL}(q_{\phi}(\boldsymbol{z}|.)\parallel p_{\phi,\omega}(\boldsymbol{z}))\right].$ (11)
We then define the overall loss of our model as
$\mathcal{L}=\mathcal{L}_{\text{rec}}-\beta\mathcal{L}_{\text{KL}}.$ (12)
This formulation is useful to learn a model that predicts the interventional distributions for a single input data set $\boldsymbol{D}^{\mathcal{M}}$, which needs to be known at training time. To successfully apply our model to data sets that were not in the training data, we expand the objective to an amortizing objective. More specifically, we consider causal models from predefined classes of models. The observational and interventional datasets are sampled from these models, inducing a distribution $p(\boldsymbol{D}^{\mathcal{M}},\boldsymbol{D}^{\mathcal{M}_{do(V)}})$ over these datasets. Optimizing our amortized objective then amounts to minimizing the expectation of our loss regarding the distribution of these datasets:
$\min_{\theta,\phi,\omega}\quad\mathbb{E}_{p(\boldsymbol{D}^{\mathcal{M}},\boldsymbol{D}^{\mathcal{M}_{do(V)}})}\mathcal{L}.$ (13)
This objective ensures that our model can make inferences on datasets coming from the class of predefined causal models, even if the specific dataset was not in the training set. It furthermore implies that the more general the training distribution of data sets is, the more the model will generalize to new data sets during inference.
@@ 4.1 Datasets
We evaluate the performance of the proposed method across four different types of datasets: two purely synthetic and one semi-synthetic dataset. Below we provide an overview of each dataset category. For each causal model $\boldsymbol{D}^{\mathcal{M}}$ in these categories, we draw $N$ samples from the observational and each interventional distribution. Detailed information on dataset generation and the exact parameters can be found in Appendix A.
@@
We generate data from linear additive causal models of the form
$V_{j}=\sum_{i\in\mathrm{Pa}(j)}\beta_{ij}\,V_{i}+\varepsilon_{j},$
where $\varepsilon_{j}$ is drawn either from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ or a Beta distribution $\mathrm{Beta}(\alpha,\beta)$. We generate data for single-variable interventions $do(V_{i}=5)$ on each variable. In general, linear Gaussian-noise models are not identifiable without additional constraints or interventional data, whereas beta-noise models are identifiable from merely observational data (Peters et al., 2017). While it is unclear whether this result transfers to the amortized learning setting in this paper, this aspect makes these two classes of data interesting for comparison.
@@
We generate biologically inspired data using the SERGIO simulator (Dibaeinia & Sinha, 2020) for gene-expression. SERGIO models single-cell gene expressions with the resulting data aligning closely with real gene-expression patterns. Notably, we use the implementation provided by (Lorch et al., 2022) that allows interventions in the simulator. We generate data for single-variable interventions $do(V_{i}=0)$ on each variable to simulate gene-knockout.
4.2 Metrics
To assess the fidelity of the learned distributions, we rely on three sample-based metrics and a permutation-based statistical test. A detailed introduction to these metrics can be found in Appendix B.
Maximum Mean Discrepancy (MMD).
MMD (Gretton et al., 2012) is a kernel-based measure that compares the mean embeddings of two distributions in a reproducing kernel Hilbert space. Smaller values indicate a closer match in the induced feature space.
Wasserstein Distance (WSD).
WSD (Villani, 2009), captures the minimal “transport cost” required to transform one distribution into another, providing a geometric notion of dissimilarity. Lower values indicate closeness of the samples.
Energy Distance (ERG).
The energy distance (Székely & Rizzo, 2013) indicates pairwise differences between samples from two distributions. A lower energy distance corresponds to higher distributional similarity.
Energy-Based Permutation Test.
We further employ a permutation test using the energy distance to evaluate whether the estimated samples differ significantly from ground-truth data. By comparing the observed energy distance against a null distribution formed via random permutations of the pooled samples, we obtain a statistical measure of distributional mismatch. The higher the resulting p-value, the more our samples are indistinguishable from the samples drawn from the ground-truth distribution.
4.3 Conditional Baseline
We compare with a baseline that estimates the interventional distribution $p(\boldsymbol{V}\mid\mathrm{do}(\boldsymbol{V}_{\boldsymbol{t}}=v_{i}))$ using conditional distributions. Concretely, it fits a multivariate normal (MVN) to the dataset in order to approximate the observational distribution $p(\boldsymbol{V})$. From the fitted MVN, it derives the conditional distribution $p(\boldsymbol{V}\setminus\boldsymbol{V}_{\boldsymbol{t}}\mid\boldsymbol{V}_{\boldsymbol{t}}=v_{i})$. Then, it draws $N$ samples from this conditional distribution and assigns the intervention value $v_{i}$ to each variable in $\boldsymbol{V}_{\boldsymbol{t}}$. This approach provides a straightforward conditional baseline, enabling us to compare the degree to which our method captures genuinely causal information, rather than relying solely on conditioning.
4.4 Demonstration of Capturing the Underlying Causal Model
In this experiment, we examine whether our model’s inference procedure indeed captures the causal model of the data-generating process. Specifically, we train our model on the Beta and Gaussian datasets, both described in Section 4.1. We first consider the bivariate case where each dataset consists of pairs of variables $(X,Y)$ under distinct causal mechanisms. Next, we extend this evaluation by training ACTIVA on Gaussian and Beta datasets with 8 variables to show its ability to scale to more complex scenarios. In particular, we train two more models for each of the two datasets. One with a single component (Gaussian1 and Beta1) and one with a mixture of 10 components, (Gaussian10 and Beta10), for the output distribution. This facilitates better comparison to the baseline since it only has one Gaussian for the output distribution as well, while demonstrating the full modeling power of our model. Details about hyperparameter settings are presented in Appendix C.
Quantitative Evaluation.
To assess our method’s performance, we compare it against the conditional baseline on a test set of novel causal models. Table 1 reports the average scores on the two test sets in the bivariate case. Firstly, we observe that ACTIVA outperforms the baseline across all metrics. Secondly, we note that for both datasets, the estimated interventional distributions of our method are not significantly different from the ground-truth interventional distributions on average [^3] Using a 5% significance level., whereas the conditional baseline’s estimates are significantly different. This demonstrates that the distributions that are estimated with ACTIVA are indistinguishable from the ground-truth distribution on average.
Table 1: Average inference performance of our trained model and the baseline on held-out test data with two variables.
MMD WSD ERG P
Gaussian baseline .92 6.06 11.13 .00
ACTIVA .36 2.79 2.57 .06
Beta baseline .91 6.02 11.45 .00
ACTIVA .30 2.66 2.52 .20
Table 2 shows the inference performance on the Gaussian test sets and the conditional baseline for the eight variables case. Results for the Beta models are presented in  Section D.1. Even in this more complex task, we observe that ACTIVA outperforms the conditional baseline in terms of MMD and ERG. This becomes even more apparent when we increase the expressiveness of the predicted distribution by adding more components to the Gaussian mixture. In this case, our model’s estimates are almost indistinguishable from the ground truth distribution with an average p-value of $0.048$. In contrast, the higher WSD reflects a greater “transport cost” in the distribution space, which we attribute to ACTIVA’s tendency to predict higher variances on average. [^4] Average variance of 0.93 for the baseline, 38.8 and 35.91 for Gaussian1 and Gaussian10, respectively.
Table 2: Average inference performance of our trained models and the baseline on held-out test data with eight variables. Gaussian1 and Gaussian10 are the models trained on the Gaussian data having single component and mixture of $10$ components, respectively.
mmd wsd erg p
Baseline .89 10.62 16.13 .00
Gaussian 1 .54 13.93 11.12 .001
Gaussian 10 .49 11.83 9.31 .048
(a) Ground Truth
@@ 4.4 Demonstration of Capturing the Underlying Causal Model
Figure 3: Samples from the observational and interventional distributions of the ground-truth model (a), the conditional baseline (b) and our model (c) for the Beta data. The observational distribution (turquoise) of the original data and the conditional baseline each have a mean of $(0.5,0.49)$. The distribution $p(Y\mid do(X=5))$ (blue) for the original data, the baseline and our model has a mean of $0.54$, $1.33$, and $0.46$, respectively. The distribution $p(X\mid do(Y=5))$ (brown) has a mean of $0.48$, $1.28$, and $0.46$, respectively. Note that there are no samples from the observational distribution from our model, since it does not estimate observational distributions. Instead we provide the observational samples of the original data for visualization.

### v2

@@ 5 Model Architecture
Our model employs a transformer-based encoder-decoder architecture that estimates interventional distributions from observational data and intervention queries. For detailed hyperparameters, see Appendix C.
@@ 5.1 Representing Interventions
To represent interventions, we create a matrix representation of the interventions. We consider $i\in\{1,\dots,|I|\}$ the index of a possible intervention value $v_{i}$, where $|I|$ is the number of possible values, and a selector $\boldsymbol{t}\in\{0,1\}^{d}$ indicating the target(s) of the intervention. By not encoding the intervention value directly, but rather the index of the intervention, we remove the need to have a full specification of the intervention. This allows us to model interventional distributions based on placeholder IDs of the interventions, as long as they can be attributed to interventional training data and have specified targets.
@@
We construct the intervention representation $\boldsymbol{I}_{i,t}$ representing $do(\boldsymbol{V}_{\boldsymbol{t}}=v_{i})$ as follows. We perform a one-hot encoding of $i$ creating a vector $\boldsymbol{i_{oh}}\in\mathbb{R}^{\mid I\mid}$ and repeat it $d$ times, resulting in a matrix $\boldsymbol{i_{rep}}\in\mathbb{R}^{d\times\mid I\mid}$. We then apply $\boldsymbol{t}$ as a mask to this matrix, effectively zeroing out the rows that correspond to non-intervened variables. Finally, we repeat the intervention representation $N$ times to match the number of input samples and obtain $\boldsymbol{I}_{i,t}\in\mathbb{R}^{N\times d\times\mid I\mid}$.
@@
This construction ensures that the intervention-relevant information for each variable is provided as local information alongside the variable itself and maintains permutation equivariance.
5.2 Embedding Network
@@
We encode the input data and intervention according to an embedding network $h_{\alpha}(\boldsymbol{D}^{\mathcal{M}},\boldsymbol{I}_{i,t})$ based on the extension of non-parametric encoders (Kossen et al. 2021; Lorch et al. 2022). For various causal tasks, this architecture can successfully encode a dataset into a vector containing causally relevant information, such as structure (Lorch et al. 2022), topological ordering (Scetbon et al. 2024), informative interventions (Annadani et al. 2025), and effects (Balazadeh et al. 2025; Robertson et al. 2025; Ma et al. 2025).
@@
Given a dataset $\boldsymbol{D}^{\mathcal{M}}\in\mathbb{R}^{N\times d}$ of $N$ observational samples from a model $\mathcal{M}$, an intervention index $i$ and a binary target vector $\boldsymbol{t}$, we append $\boldsymbol{I}_{i,t}$ to the dataset resulting in an augmented dataset $\boldsymbol{D}^{\mathcal{M}}_{it}\in\mathbb{R}^{N\times d\times\mid I\mid+1}$. Then we apply $L$ blocks of multi-head self-attention (MHSA) that alternate in attending over $d$ features and $N$ samples. Furthermore, we apply layer normalization, residual connections, and dropout following the transformer setup (Vaswani et al. 2017). After the transformer blocks, we average over the sample axis to obtain an embedding $\boldsymbol{h}\in\mathbb{R}^{d\times e}$, where $e$ indicates the embedding dimension, thus ensuring that the embedding is permutation invariant regarding the sample dimension and equivariant regarding the feature dimension.
5.3 Prior and Encoder
Following the reasoning in Section 4.1, we choose the conditional prior to be an n-dimensional GMM with constant variance and no covariance whose means, and weights are determined by a learnable prior function $\boldsymbol{\mu}_{0:d},\rho_{0:C}]=pri_{\eta}(h_{\alpha^{\prime}}(\boldsymbol{D}^{\mathcal{M}},\boldsymbol{I}_{i,t}))$. Specifically, $pri_{\eta}$ are two two-layer multi-layer perceptron (MLP) that we apply to the embedding $\boldsymbol{h}^{\prime}$ to each feature independently to compute the means and mixture weights. The outputs parametrize our prior distribution $p_{\eta}(\boldsymbol{z}|\boldsymbol{D}^{\mathcal{M}},do(V))=\sum_{j=0}^{|C|}\rho_{j}\mathcal{MVN}(\boldsymbol{z};\boldsymbol{\mu_{j}},0.1^{C})$, with and $\boldsymbol{\mu}_{j}$ the mean vector of each component, respectively. We enforce $\sum_{j=0}^{C}\rho_{j}=1$ via a softmax layer.
Furthermore, we choose the encoder distribution to be an n-dimensional MVN with constant, diagonal covariance matrix whose means are determined by a learnable encoder function $\boldsymbol{\mu}=enc_{\phi}(h_{\alpha^{\prime\prime}}(\boldsymbol{V},\boldsymbol{I}_{i,t}),h_{\alpha^{\prime}}(\boldsymbol{D}^{\mathcal{M}},\boldsymbol{I}_{i,t}))$, where $h_{\alpha^{\prime}}$ is the observational data network used in the prior, and $h_{\alpha^{\prime\prime}}$ is the network encoding interventional data with different parameters. We concatenate the resulting embeddings $\boldsymbol{h}^{\prime}$ and $\boldsymbol{h}^{\prime\prime}$ along the embedding dimensions and again apply a two-layer MLP to each feature independently to compute the mean. This fully parametrized our encoding distribution $q_{\phi}(\boldsymbol{z}|\boldsymbol{V},\boldsymbol{D}^{\mathcal{M}},do(V))=\mathcal{MVN}(\boldsymbol{z};\boldsymbol{\mu},0.1)$.
@@ 5.4 Decoder
We model the decoder $dec_{\gamma}(\boldsymbol{z})$ as a transformer that outputs the parameters of a Gaussian mixture, enabling a closed-form expression for the estimated interventional distributions.
For the decoder we simply process $\boldsymbol{z}$ via $K$ standard transformer blocks (Vaswani et al. 2017) resulting in the embedding $\boldsymbol{h}_{dec}$. We feed this embedding through a row-wise linear layer to predict the means $\boldsymbol{\mu}_{\gamma}\in\mathbb{R}^{B\times d}$, where $B$ is the number of decoder mixture components, maintaining permutation equivariance regarding the variable ordering. To also predict the covariances $\boldsymbol{\Sigma}_{\gamma}\in\mathbb{R}^{B\times d\times d}$ in a permutation equivariant manner, we first apply a row-wise linear layer to compute $\boldsymbol{u}\in\mathbb{R}^{B\times d\times e}$ as an intermediate step and then compute the covariances via $\boldsymbol{u}\cdot\boldsymbol{u}$ similarly to (Lorch et al. 2022). Lastly, we predict the component weights $\boldsymbol{b}_{0:B}$ of the mixture by first summing $\boldsymbol{h}_{dec}$ along the $d$ feature dimensions and then passing the pooled representation through a linear layer and a softmax layer.
We note that we maintain permutation equivariance of the means and covariances of the resulting mixture w.r.t. the ordering of the variables end-to-end, a property that has been shown to be important in causal inference for scaling (Lorch et al. 2022) and accurracy (Li, Xiao, and Tian 2020), among others.
@@ 6.1 Datasets
We evaluate the performance of the proposed method across three different types of datasets: two purely synthetic and one semi-synthetic dataset. Below we provide an overview of each dataset category. Detailed information on dataset generation and the exact parameters can be found in A.
@@
We generate data from linear additive causal models of the form $V_{j}=\sum_{i\in\mathrm{Pa}(j)}\beta_{ij}\,V_{i}+\varepsilon_{j},$ where $\varepsilon_{j}$ is drawn either from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ or a Beta distribution $\mathrm{Beta}(\alpha,\beta)$. We generate data for single-variable interventions $do(V_{i}=5)$ on each variable. In general, linear Gaussian-noise models are not identifiable without additional constraints or interventional data, whereas beta-noise models are identifiable from merely observational data (Peters, Janzig, and Schölkopf 2017), which makes these two model classes interesting for comparison.
@@
We generate biologically inspired data using the SERGIO simulator (Dibaeinia and Sinha 2020) for gene-expression. SERGIO models single-cell gene expressions with the resulting data aligning closely with real gene-expression patterns. Notably, we use the implementation provided by (Lorch et al. 2022) that allows interventions in the simulator. We generate data for single-variable interventions $do(V_{i}=0)$ on each variable to simulate gene-knockout.
(a) Ground-truth
@@
Figure 2: Samples from the observational and interventional distributions of the ground-truth model (a), the conditional baseline (b) and our model (c) for the Beta data. The turquoise represent samples from the observational distribution, blue and brown depict samples from the interventional distributions by setting $X$ and $Y$ to 5, respectively. Note that for visualization purposes, we provide the observational samples of the original data for the ACTIVA plot.
6.2 Metrics
To assess the fidelity of the learned distributions, we rely on the following three sample-based metrics and a permutation-based statistical test to get a broad picture of the distributional accuracy of our estimations. For all three metrics, lower values are desired. A detailed introduction to these metrics can be found in Appendix B.
Maximum Mean Discrepancy (MMD) (Gretton et al. 2012) is a kernel-based measure that compares the mean embeddings of two distributions.Wasserstein Distance (WSD) (Villani 2009), captures the minimal “transport cost” required to transform one distribution into another. The energy distance (ERG) (Székely and Rizzo 2013) indicates pairwise differences between samples.
6.3 Conditional Baseline
We compare with a baseline that estimates the interventional distribution $p(\boldsymbol{V}\mid\mathrm{do}(\boldsymbol{V}_{\boldsymbol{t}}=v_{i}))$ using conditional distributions. Concretely, it fits a multivariate normal (MVN) to the dataset in order to approximate the observational distribution $p(\boldsymbol{V})$. From the fitted MVN, it derives the conditional distribution $p(\boldsymbol{V}\setminus\boldsymbol{V}_{\boldsymbol{t}}\mid\boldsymbol{V}_{\boldsymbol{t}}=v_{i})$. Then, it draws $N$ samples from this conditional distribution and assigns the intervention value $v_{i}$ to each variable in $\boldsymbol{V}_{\boldsymbol{t}}$. This approach provides a baseline, enabling us to compare the degree to which our method captures genuinely causal information, rather than relying solely on conditioning.
ACTIVA Baseline
Data MMD WSD ERG MMD WSD ERG
Gauss 2 0.29 1.9 2.0 0.91 8.0 7.9
Beta 2 0.21 1.4 1.4 0.90 5.8 11.0
Gauss 8 0.65 11.4 12.9 0.88 19.9 16.4
Beta 8 0.77 9.3 13.7 0.94 11.2 19.0
Table 1: Average inference performance of our trained model and the baseline on held-out test data with 2 and 8 variables respectively. The Data column indicates which dataset has been used with how many variables.
6.4 Distribution Shift Estimation
In this experiment, we examine whether our model’s inference procedure indeed captures causal distribution shifts. Specifically, we train our model on the Beta and Gaussian datasets, both described in Section 6.1. We first consider the bivariate case where each dataset consists of pairs of variables $(X,Y)$ under distinct causal mechanisms. Next, we extend this evaluation by training ACTIVA on Gaussian and Beta datasets with 8 variables to show its ability to scale to more complex scenarios. Details about hyperparameter settings are presented in Appendix C.
Quantitative Analysis.
We assess our method’s performance, by comparing it against the conditional baseline on a test set of novel causal models. The models are novel in the sense that while they come from the same distribution of models $p_{tr}(\mathcal{M})$, the models that were sampled in the test set are not in the training set. Table 1 reports the average scores on the two test sets in the bivariate and 8 variable case performing inference with one sample from the prior.
We observe that ACTIVA outperforms the baseline across all metrics and for both model sizes. This validates our implementation of ACTIVA to be able to recover causal distribution shifts beyond conditioning. Furthermore, we observe in the 2 variable case the increased performance on the Beta data compared to the Gaussian data. We interpret this as reflecting the identifiability of the Beta data compared to the general non-identifiability of the Gaussian data. Lastly, we attribute the comparably lower performance margin of the 8 variable models w.r.t. the baseline to non-identifiability in our data (see Assumption 1), as we only provide data with one-variable interventions during training.